40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2559     Entry time: Thu Mar 12 19:11:00 2020
Author: shruti 
Type: DailyProgress 
Category: ISS 
Subject: Removed half-wave plate in north path 

[Anchal, Shruti]

We realized that the half-wave plates before the EOAMs probably had no real function in the setup and therefore we proceeded to remove the one from the north path at (39,121) aka row 39, column 121 of the Optical layout.

After this was done, we had to re-adjust the quarter wave-plate (39,112) after the EOAM (39,115) to make sure that the EOAM was still functioning about the 50% transmission point. The beam going into the PMC was also re-aligned by adjusting the two mirrors at (32,92) and (37,92). Finally, the mirror at (43,88) was adjusted to align the beam reflecting from the PMC into the photo-diode.

We were able to re-lock the north PMC and north cavity after increasing the power in that path by adjusting some waveplates.

As may be expected, the sign of the ISS feedback had to be inverted. The ISS actuates on the EOAM; removing the half-wave plate would have switched the circularity of the polarization of the beam entering the PBS at (39,110), so the sign of the voltage that would have previously caused the transmission to increase would now cause it to decrease and vice versa.

ELOG V3.1.3-